You read in BusinessWeek that a panel of economists has estimated that the long-run real growth rate of the U.S. economy over the next five-year period will average 8 percent. In addition, a bank newsletter estimates that the average annual rate of inflation during this five-year period will be about 5 percent. What nominal rate of return would you expect on U.S. government T-bills during this period? Round your answer to two decimal places.

  %

What would your required rate of return be on common stocks if you wanted a 6 percent risk premium to own common stocks? Do not round intemediate calculations. Round your answer to two decimal places.

  %

If common stock investors became more risk averse, what would happen to the required rate of return on common stocks? What would be the impact on stock prices?

As an investor becomes more risk averse, the required rate of return will _____ 3 and the stock prices will _____
Answer #1

Answers:-

1) Nominal rate of return =(1+growth rate)(1+annual rate) -1

=(1+8%)(1+5%) -1 =0.134 or 13.4%

(Approximation=8% + 5% = 13%)

2)Return on common stock =(1+ nominal rate)(1+rate of return) -1

=(1+13.4%)(1+6%) -1 = 0.20204 or 20.20%

(Approximation = 8%+ 5% +6% = 19%)

3) As an investor becomes more risk averse , the investor will require a larger risk premium to own common stock.

As risk premium increases, so too will required rate of return.

In order to acheive the high rate of return ,stock prices should be decline.